OpenStudy (anonymous):

Equation 1 : 5x-2x = 2x + 2 Equation 2 : 5x+2=2x-10 A the solution to equation #1is smaller than the solution to equation #2 B the solution to equation #1 is larger than the solution to equation #2 C the solution to equation #1 is the same as the solution to equation #2 D none of the statements above describe solutions to the equations shown I WILL FOLLOW AND GIVE MEDAL

OpenStudy (anonymous):

I will follow annd give a medal

OpenStudy (anonymous):

A

OpenStudy (anonymous):

The first one totals to two and the second totals to four

OpenStudy (anonymous):

5x + -2 = 2x + 10 Reorder the terms: -2 + 5x = 2x + 10 Reorder the terms: -2 + 5x = 10 + 2x Solving -2 + 5x = 10 + 2x Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-2x' to each side of the equation. -2 + 5x + -2x = 10 + 2x + -2x Combine like terms: 5x + -2x = 3x -2 + 3x = 10 + 2x + -2x Combine like terms: 2x + -2x = 0 -2 + 3x = 10 + 0 -2 + 3x = 10 Add '2' to each side of the equation. -2 + 2 + 3x = 10 + 2 Combine like terms: -2 + 2 = 0 0 + 3x = 10 + 2 3x = 10 + 2 Combine like terms: 10 + 2 = 12 3x = 12 Divide each side by '3'. x = 4 Simplifying x = 4

OpenStudy (anonymous):

Its a or b

OpenStudy (anonymous):

Thx

OpenStudy (anonymous):

Can you help me with 1 more